Problem 7 (20 points) Innovation Company is thinking about marketing a new software product Up-front costs to market and develop the product are $5 million. The product is expected to generate profits of $1 million per year for 10 years. The company will have to provide product support expected to cost $100,000 per year in perpetuity. Assume all profits and expenses occur at the end of the year. what is the NPV of this investment if the cost of capital is 6%? Should the firm undertake the project? (10 points) Repeat the analysis for discount rates of 2% and 11%. (10 points) a. b.
Answer #1

Cost to develop and market = $5 million

At 6% cost of capital,

Maintenance cost to perpetuity = 0.1/0.06 = $1.67 million

PV =-I t + XC,/ (1 + r)-Costof Maintenance natial Investmen

NPV = $2.36 - 1.67

NPV = $0.69 million

At 2% cost of capital,

Maintenance cost to perpetuity = 0.1/0.02 = $5 million

PV =-I t + XC,/ (1 + r)-Costof Maintenance natial Investmen

NPV = 3.98 - 5

NPV = -$1.02 million

At 11% cost of capital,

Maintenance cost to perpetuity = 0.1/0.11 = $0.91 million

PV =-I t + XC,/ (1 + r)-Costof Maintenance natial Investmen

NPV = 0.89 - 0.91

NPV = -$0.02 million
Answer #2
Multiple Internal Rates of Return: Problem 6-13 in Berk   and DeMarzo (2010)
Cost of   capital6%Data Table: NPV of the   project as a function of the discount rate
upfront   costs -$5.00


Annual   profits$1.00
$0.69
Annual   support-$0.101%-$5.53
Production   period, Years102%-$1.02


3%$0.20
PV   support-$1.674%$0.61
PV of   profit before depreciation$7.365%$0.72
NPV$0.696%$0.69


7%$0.60


8%$0.46


9%$0.31


10%$0.14


11%-$0.02


12%-$0.18


13%-$0.34


14%-$0.50


15%-$0.65


16%-$0.79


17%-$0.93


18%-$1.06


19%-$1.19


20%-$1.31


answered by: Sheetal Kdinesh